OpenStudy (anonymous):

A bike travles 12.5 miles in 2 hours and 30 minutes and 27.5 miles in 5 hours and 30 minutes. What is the constant of proportionally?

OpenStudy (ranga):

Distance = k * time 12.5 = k * 2.5 k = 12.5/2.5 = 5 Make sue k works for the other data: 27.5 miles in 5.5 hours. distance = 5 * time put time = 5.5 distance = 5 * 5.5 = 27.5 miles which matches the data given in the problem. So k = 5 miles/hr.